OpenStudy (anonymous):

The Good Seed Flower Shop sells its own brand of fertilizer. It mixes one brand of fertilizer that contains 7% nitrogen with another than contains 8% nitrogen. The resulting mixture contains 7.7% nitrogen. How many pounds of the 8% fertilizer must be used to make 200 pounds of the 7.7% mixture?

OpenStudy (whpalmer4):

Let's call the variable representing the 7% nitrogen \(s\) Let's call the variable representing the 8% nitrogen \(e\) With mixture problems, it is often advantageous to work in terms of the amount of the thing of interest. Here, that would be the nitrogen. \(0.07s\) gives the number of pounds of nitrogen in \(s\) pounds of 7% fertilizer \(0.08e\) gives the number of pounds of nitrogen in \(e\) pounds of 8% fertilizer We are going to mix \(s\) pounds of 7% with \(e\) pounds of 8% to make 200 pounds of the 7.7% mixture, so we can write\[s+e=200\] The amount of nitrogen in our mixture is going to be the amount of nitrogen in 200 pounds of 7.7% mixture, so we can also write \[0.07s+0.08e = 200*0.077\] Now you have a system of two equations in two unknowns, which can be easily solved by substitution.